# Kinetic Theory of Quasiparticles, Retarded Correlators and Hydrodynamics

### v1

@@ IV.1 Hydrodynamic limit
After obtaining the correlation functions, we must match their small momentum and frequency limits to the universal hydrodynamic behaviors. The hydrodynamic limit of the no-pole correlation function $G^{12,12}(\omega,\vec{k})$ for a non-conformal system up to the third-order transports is known as follows [56]
$G^{12,12}(\omega,\vec{k})$ $=P_{eq}-i\eta\omega+\left(\eta\tau_{\pi}-\frac{\kappa}{2}+\kappa^{*}\right)\omega^{2}-\frac{\kappa}{2}k^{2}$
@@
To obtain the transport coefficient, we should derive the expression of $G^{12,12}(\omega,\vec{k})$ provided in Eq. (IV). The low $x$ expansion of this correlation function takes the following form
$G^{12,12}(\omega,\vec{k})=i\tau_{eq}\omega\,\frac{\varepsilon_{eq}+P_{eq}}{8c_{s}^{2}}\bigg{\{}\frac{(1-i\tau\omega)^{3}}{(k\tau)^{4}}+\frac{5(1-i\tau\omega)c_{s}^{2}}{(k\tau)^{2}}+\frac{(1-i\tau\omega)\left(k^{2}\tau^{2}+(1-i\tau\omega)^{2}\right)g_{1}(x)}{(k\tau)^{4}}$
$\hskip 56.9055pt-\frac{(1-i\tau\omega)\left(7k^{2}\tau^{2}+(1-i\tau\omega)^{2}\right)g_{2}(x)}{(k\tau)^{4}}+\frac{i}{2}L(\omega,\vec{k})\mathcal{G}(\omega,k,x)\bigg{\}},$ (51)
@@
We perform the small $(\omega,k)$ expansion of Eq. (IV.1) and compare it to Eq. (IV.1). This shall give us transport coefficients
$\frac{\eta}{T_{0}^{3}}=\frac{xT_{0}\tau_{eq}}{30\pi^{2}}\bigg{(}x(12-x^{2}-3x^{4})K_{0}(x)+(24+x^{2}+x^{4})K_{1}(x)\bigg{)},$
$\frac{\eta}{S_{eq}}=\frac{T_{0}\tau_{eq}}{15x^{2}K_{3}(x)}\bigg{(}x(12-x^{2}-3x^{4})K_{0}(x)+(24+x^{2}+x^{4})K_{1}(x)\bigg{)},$
$\kappa=\kappa^{*}=0,$
$\tau_{\pi}=\tau_{eq},$
$\upsilon^{(3,2)}_{31}+\upsilon^{(3,2)}_{32}-\upsilon^{(3,2)}_{34}=2\eta\tau^{2}_{eq},$
$\upsilon^{(3,2)}_{1}-\upsilon^{(3,2)}_{31}-\upsilon^{(3,2)}_{32}=\frac{xT_{0}^{4}\tau_{eq}^{3}}{105\pi^{2}}\bigg{(}2x(6-x^{2}-3x^{4})K_{0}(x)+(24-x^{2}+3x^{4})K_{1}(x)\bigg{)},$ (53)
We can obtain the small $x$ expansion and compare it to the massless case [27]
@@
It is worth mentioning that the constraint $\eta\geq 0$ would imply that $x\leq 1.847$. This limit is larger than the criteria $x\leq 1.2342$ derived in the previous section to ensure stability. It means that the stability of diffusive modes will respect the second law of thermodynamics.
@@
$c_{0}=1-\frac{2c_{s}^{2}}{i\tau_{eq}\omega(\varepsilon_{eq}+P_{eq})}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{4}(v,\omega,\vec{k})-\frac{c_{s}^{2}(1-i\tau_{eq}\omega)}{k^{2}\tau_{eq}^{2}(\varepsilon_{eq}+P_{eq})}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{1}(v,\omega,\vec{k}),$
$c_{3}=-\frac{2}{\varepsilon_{eq}+P_{eq}}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{6}(v,\omega,\vec{k})-\frac{1}{\varepsilon_{eq}+P_{eq}}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{5}(v,\omega,\vec{k}).$ (58)
@@
To respect the second law of thermodynamics, which is manifested as $\xi\geq 0$, we constrain to $x\leq 1.62$. In Fig. 3, we plot $\xi/\eta$ in terms of $x$. The maximum value of $\xi/\eta$ occurs at $x\sim 1.452$.
@@
In what follows, we will derive the hydro modes in the sound channel. To do this, we expand the $G^{00,00}$ for small $(\omega,k)$ that leads to the following expression
$G^{00,00}(\omega,k)\bigg{|}_{\omega,k\ll 1}=\frac{(\varepsilon_{eq}+P_{eq})\,i\,k^{2}\tau_{eq}\bigg{(}-1+2g_{1}(x)-8g_{2}(x)\bigg{)}}{\omega\bigg{(}1-3c_{s}^{2}(1-i\tau_{eq}\omega)-2g_{1}(x)+8g_{2}(x)\bigg{)}\,\mathcal{D}_{sound}(\omega,k,x)},$ (61)
@@
$\mathcal{D}_{sound}(\omega,k,x)=1+\frac{c_{s}^{2}\,k^{2}\,\tau_{eq}\mathcal{F}_{sound}^{(1)}(\omega,x)}{\mathcal{F}_{sound}^{(2)}(\omega,x)},$ (62)
@@
$\omega_{1}=\frac{ic_{s}^{2}k^{2}\tau_{eq}(2g_{1}(x)-8g_{2}(x)-1)}{-1+3c_{s}^{2}+2g_{1}(x)-8g_{2}(x)}+\mathcal{O}(k^{4}),$ (63)
$\omega_{2}=-\frac{i}{\tau_{eq}}-\frac{4ic_{s}^{2}k^{2}\tau_{eq}\,\bigg{(}g_{1}(x)(4-40g_{2}(x))+5g_{1}(x)^{2}+80g_{2}(x)^{2}-34g_{2}(x)-1\bigg{)}}{5\left(1-2g_{1}(x)+8g_{2}(x)\right)^{2}}+\mathcal{O}(k^{4}),$
$\omega_{3}=\frac{i\left(1-3c_{s}^{2}-2g_{1}(x)+8g_{2}(x)\right)}{3c_{s}^{2}\tau}$
$+\frac{3}{5}ic_{s}^{2}k^{2}\tau_{eq}\bigg{(}\frac{5c_{s}^{2}}{3c_{s}^{2}+2g_{1}(x)-8g_{2}(x)-1}-\frac{9}{1-2g_{1}(x)+8g_{2}(x)}-\frac{6(g_{1}(x)-1)}{(1-2g_{1}(x)+8g_{2}(x))^{2}}\bigg{)}+\mathcal{O}(k^{4}).$
The massless limit of the pole $\omega_{2}$ continuously merge into the formal result $\omega=-\frac{i}{\tau_{eq}}+\frac{4ik^{2}\tau_{eq}}{15}$, but the massless limit of other two poles, namely $\omega_{1}$ and $\omega_{3}$ is different from analogous result given in [27]. Indeed, the mass disappears the propagating sound modes and turns them into dissipative modes.
@@
$G^{0i,0i}(\omega,k)\bigg{|}_{\omega,k\ll 1}=\frac{c_{s}^{2}\,\bigg{(}70(1-i\tau_{eq}\omega)^{2}\,\left(-1+3c_{s}^{2}+2g_{1}(x)-8g_{2}(x)\right)+k^{2}\tau_{eq}^{2}\mathcal{N}_{shear}(x)\bigg{)}}{70(1-i\tau_{eq}\omega)\,\bigg{(}1-3c_{s}^{2}(1-i\tau_{eq}\omega)\,-2g_{1}(x)+8g_{2}(x)\bigg{)}\bigg{(}1-\frac{k^{2}\tau_{eq}^{2}\mathcal{D}_{shear}(\omega,x)}{6160(1-i\tau_{eq}\omega)^{2}}\bigg{)}},$ (64)
@@
$\mathcal{N}_{shear}(x)=\frac{32-180c_{s}^{2}+8g_{1}(x)\left(135c_{s}^{2}-93g_{2}(x)-32\right)-3g_{2}(x)\left(780c_{s}^{2}+461g_{2}(x)-112\right)+272g_{1}(x)^{2}}{8g_{1}(x)-33g_{2}(x)-2},$
$\mathcal{D}_{shear}(\omega,x)=461+\frac{77\left(99c_{s}^{2}(1-i\tau_{eq}\omega)+2g_{1}(x)-17\right)}{1-3c_{s}^{2}(1-i\tau_{eq}\omega)-2g_{1}(x)+8g_{2}(x)}+\frac{160(7-94g_{1}(x))}{8g_{1}(x)-33g_{2}(x)-2}.$ (65)
@@
$\omega_{1,2}=-\frac{i}{\tau_{eq}}\pm k\sqrt{\frac{32-8g_{1}(x)(93g_{2}(x)+32)+272g_{1}(x)^{2}+3g_{2}(x)(112-461g_{2}(x))}{70\left(8g_{1}(x)-33g_{2}(x)-2\right)\left(2g_{1}(x)-8g_{2}(x)-1\right)}}+\mathcal{O}(k^{2}),$
$\omega_{3}=\frac{i\left(1-3c_{s}^{2}-2g_{1}(x)+8g_{2}(x)\right)}{3c_{s}^{2}\tau_{eq}}+\frac{3ic_{s}^{2}k^{2}\tau_{eq}(8g_{1}(x)-33g_{2}(x)-2)}{10(1-2g_{1}(x)+8g_{2}(x))^{2}}+\mathcal{O}(k^{4}).$ (66)
The massless limit of the latter equation coincides with the analogous results in [27].
@@ V Analytical Behaviors
$\frac{1}{\omega}\mbox{Im}\,G_{00}^{(l)}(x;\omega,k=0)\bigg{|}_{\tau_{eq}T\to\infty}=\pi\chi(x)\,\delta(\omega)\,\left(1+\frac{x^{2}K_{0}(x)+4xK_{1}(x)}{8K_{2}(x)}\right),$ (68)
which is always positive and respects the positivity condition of the spectral density. In the strong coupling regime, the expansion of the retarded two-point function matches Eq. (28), demonstrating that hydrodynamic poles emerge in this regime.
@@
This solution corresponds to the small-mass limit of $\omega_{3}^{(h)}$ in Eq. (III.1) and coincides at $x=0$ with the results in [27]. In Fig. 4, we show the lowest-lying poles of the retarded $G_{00}$ correlator given in Eq. (28) for different $x$ values in the strong coupling regime at various $k\tau_{eq}$. Dashed lines indicate the positions of the branch cuts, and the hydrodynamic poles of the massless system are marked with crosses. Similar to Fig. 2, the mass tends to shift the poles toward the unstable region.
@@
$\frac{-1}{\omega}\mbox{Im}\,G^{00,00}(\omega,k=0)\bigg{|}_{\tau_{eq}T\to\infty}=\frac{\varepsilon_{eq}+P_{eq}}{c_{s}^{2}}\,\pi\,\delta(\omega)\,\left(1+g_{1}(x)-g_{2}(x)\right),$ (71)
which satisfies the positive spectral density condition. In the strong coupling limit, the correlator expansion coincides with Eq. (61), and hydrodynamic poles naturally emerge. In Fig. 5, we depict the lowest-lying poles of $G^{00,00}$ for different $x$ values in the strong coupling regime at various $k\tau$. Unlike the massless case, all poles are located on the vertical line in the unstable region. These numerical solutions agree with those given in Eq. (63).
Figure 4: Comparison between hydro poles for massless and massive cases in the charge correlation functions. In each figure, the dashed line indicates the location of the branch cut, and massless poles are shown with cross marks. Various “$x$” ratios are shown with different colors. Left (right) top plots correspond to $k\tau_{eq}=0.2(k\tau_{eq}=0.5)$, and left (right) bottom plots demonstrate $k\tau_{eq}=1(k\tau_{eq}=1.5)$.

### v2

@@ IV.1 Hydrodynamic limit
After obtaining the correlation functions, we must match their small momentum and frequency limits to the universal hydrodynamic behaviors. The hydrodynamic limit of the no-pole correlation function $G^{12,12}(\omega,k)$ for a non-conformal system up to the third-order transports is known as follows [56]
$G^{12,12}(\omega,k)$ $=P_{eq}-i\eta\omega+\left(\eta\tau_{\pi}-\frac{\kappa}{2}+\kappa^{*}\right)\omega^{2}-\frac{\kappa}{2}k^{2}$
@@
To obtain the transport coefficient, we should derive the expression of $G^{12,12}(\omega,k)$ provided in Eq. (IV). The small $x$ expansion of this correlation function takes the following form
$G^{12,12}(\omega,k)\bigg{|}_{x\ll 1}=i\tau_{eq}\omega\,\frac{\varepsilon_{eq}+P_{eq}}{8c_{s}^{2}}\bigg{\{}\frac{(1-i\tau\omega)^{3}}{(k\tau)^{4}}+\frac{5(1-i\tau\omega)c_{s}^{2}}{(k\tau)^{2}}+\frac{(1-i\tau\omega)\left(k^{2}\tau^{2}+(1-i\tau\omega)^{2}\right)g_{1}(x)}{(k\tau)^{4}}$
$\hskip 56.9055pt-\frac{(1-i\tau\omega)\left(7k^{2}\tau^{2}+(1-i\tau\omega)^{2}\right)g_{2}(x)}{(k\tau)^{4}}+\frac{i}{2}L(\omega,k)\mathcal{G}(\omega,k,x)\bigg{\}},$ (52)
@@
We perform the small $(\omega,k)$ expansion of Eq. (IV.1) and compare it to Eq. (IV.1). This shall give us transport coefficients for small $x$
@@
For large $x$, the following expression is obtained
$G^{12,12}(\omega,k)\bigg{|}_{x\gg 1}=i\tau_{eq}\omega\,\frac{\varepsilon_{eq}+P_{eq}}{8c_{s}^{2}}\bigg{\{}\frac{(1-i\tau\omega)^{3}}{(k\tau)^{4}}+\frac{5(1-i\tau\omega)c_{s}^{2}}{(k\tau)^{2}}-\frac{\mathcal{H}(\omega,k,x)}{k^{4}\tau_{eq}^{4}(1-i\tau_{eq}\omega)}\bigg{\}},$ (55)
where
$\mathcal{H}(\omega,k,x)\equiv(1-i\tau_{eq}\omega)^{4}+5k^{2}\tau_{eq}^{2}(1-i\tau_{eq}\omega)^{2}c_{s}^{2}h_{1}(x)+k^{2}\tau_{eq}^{2}\left(8k^{2}\tau_{eq}^{2}-25(1-i\tau_{eq}\omega)^{2}\right)c_{s}^{2}h_{2}(x).$ (56)
We expand the latter two-point function in small $\omega$ and $k$ and compare it with Eq. (IV.1) for large $x$. The result has become as follows
$\frac{\eta}{T_{0}^{3}}=T_{0}\tau_{eq}\frac{e^{-x}x^{3/2}}{2\sqrt{2}\pi^{3/2}},$
$\upsilon^{(3,2)}_{1}-\upsilon^{(3,2)}_{31}-\upsilon^{(3,2)}_{32}=0,$
$\kappa=\kappa^{*}=0,\qquad\tau_{\pi}=\tau_{eq},$
$\upsilon^{(3,2)}_{31}+\upsilon^{(3,2)}_{32}-\upsilon^{(3,2)}_{34}=2\eta\tau^{2}_{eq}.$ (57)
@@
$c_{0}=1-\frac{2c_{s}^{2}}{i\tau_{eq}\omega(\varepsilon_{eq}+P_{eq})}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{4}(v,\omega,k)-\frac{c_{s}^{2}(1-i\tau_{eq}\omega)}{k^{2}\tau_{eq}^{2}(\varepsilon_{eq}+P_{eq})}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{1}(v,\omega,k),$
$c_{3}=-\frac{2}{\varepsilon_{eq}+P_{eq}}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{6}(v,\omega,k)-\frac{1}{\varepsilon_{eq}+P_{eq}}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{5}(v,\omega,k).$ (61)
@@
The small $x$ expansion of bulk viscosity reads as follows
$\frac{\xi}{T_{0}^{3}}\bigg{|}_{x\ll 1}=\frac{T_{0}\tau_{eq}}{1620}x^{4}+\mathcal{O}(x^{6}).$ (64)
@@
In what follows, we will derive the hydro modes in the sound channel. To do this, we expand the $G^{00,00}$ for small $(\omega,k)$ that leads to the following expression for small $x$
$G^{00,00}(\omega,k)\bigg{|}_{x\ll 1}=\frac{(\varepsilon_{eq}+P_{eq})\,i\,k^{2}\tau_{eq}\bigg{(}-1+2g_{1}(x)-8g_{2}(x)\bigg{)}}{\omega\bigg{(}1-3c_{s}^{2}(1-i\tau_{eq}\omega)-2g_{1}(x)+8g_{2}(x)\bigg{)}\,\mathcal{D}^{(s)}_{sound}(\omega,k,x)},$ (65)
@@
$\mathcal{D}^{(s)}_{sound}(\omega,k,x)=1+\frac{c_{s}^{2}\,k^{2}\,\tau_{eq}\mathcal{F}_{sound}^{(1)}(\omega,x)}{\mathcal{F}_{sound}^{(2)}(\omega,x)},$ (66)
@@
$\omega_{1}=\frac{8ik^{2}\tau_{eq}}{x^{6}(\gamma_{E}-\ln{2})}+\mathcal{O}(k^{4}),$
$\omega_{2}=-\frac{i}{\tau_{eq}}-ik^{2}\tau_{eq}(\frac{4}{15}-\frac{x^{2}}{45})+\mathcal{O}(k^{4}),$
$\omega_{3}=\frac{ix^{6}(\gamma_{E}-\ln{2})}{24\tau_{eq}}+\frac{8ik^{2}\tau_{eq}}{x^{6}(\ln{2}-\gamma_{E})}+\mathcal{O}(k^{4}),$ (67)
where $\gamma_{E}=0.5772$ is the Euler number. The massless limit of the pole $\omega_{2}$ continuously merges into the known result [27], but the massless limit of the other two poles, namely $\omega_{1}$ and $\omega_{3}$, is different from the analogous result ones. Indeed, the mass disappears the propagating sound modes and turns them into dissipative modes. Since $\gamma_{E}-\ln{2}\simeq-0.115<0$, the lowest lying modes are stable against fluctuation.
For large $x$, the retarded two-point function has following expansion
$G^{00,00}(\omega,k)\bigg{|}_{x\gg 1}=\frac{(\varepsilon_{eq}+P_{eq})\,i\,k^{2}\tau_{eq}(1-i\tau_{eq}\omega)h_{1}(x)^{3}}{\mathcal{D}^{(l)}_{sound}(\omega,k,x)},$ (68)
where
$\mathcal{D}^{(l)}_{sound}(\omega,k,x)=h_{1}(x)^{3}\left(\omega-i\tau_{eq}(\omega^{2}+c_{s}^{2}k^{2})\right)-\omega(1-i\tau_{eq}\omega)^{2}h_{1}(x)^{2}-5k\tau_{eq}\omega(1-i\tau_{eq}\omega)^{2}h_{2}(x)\bigg{(}h_{1}(x)+5h_{2}(x)k\tau_{eq}\bigg{)}.$ (69)
At large $x$, the following modes exist at the lowest level
$\omega_{1}=ik^{2}\tau_{eq}(\frac{1}{5}+\frac{1}{x})+\mathcal{O}(k^{4}),$
$\omega_{2}=\frac{5i}{\tau_{eq}x}-\frac{ik^{2}\tau_{eq}}{5}+\mathcal{O}(k^{4}),$
$\omega_{3}=-\frac{i}{\tau_{eq}}-\frac{ik^{2}\tau_{eq}}{x}+\mathcal{O}(k^{4}).$ (70)
Unlike the $(\omega_{2},\omega_{3})$ modes, which are stable, the mode $\omega_{1}$ is unstable in large $x$ limit.
@@
$G^{0i,0i}(\omega,k)\bigg{|}_{x\ll 1}=\frac{c_{s}^{2}\,\bigg{(}70(1-i\tau_{eq}\omega)^{2}\,\left(-1+3c_{s}^{2}+2g_{1}(x)-8g_{2}(x)\right)+k^{2}\tau_{eq}^{2}\mathcal{N}^{(s)}_{shear}(x)\bigg{)}}{70(1-i\tau_{eq}\omega)\,\bigg{(}1-3c_{s}^{2}(1-i\tau_{eq}\omega)\,-2g_{1}(x)+8g_{2}(x)\bigg{)}\bigg{(}1-\frac{k^{2}\tau_{eq}^{2}\mathcal{D}^{(s)}_{shear}(\omega,x)}{6160(1-i\tau_{eq}\omega)^{2}}\bigg{)}},$ (71)
@@
$\mathcal{N}^{(s)}_{shear}(x)=\frac{32-180c_{s}^{2}+8g_{1}(x)\left(135c_{s}^{2}-93g_{2}(x)-32\right)-3g_{2}(x)\left(780c_{s}^{2}+461g_{2}(x)-112\right)+272g_{1}(x)^{2}}{8g_{1}(x)-33g_{2}(x)-2},$
$\mathcal{D}^{(s)}_{shear}(\omega,x)=461+\frac{77\left(99c_{s}^{2}(1-i\tau_{eq}\omega)+2g_{1}(x)-17\right)}{1-3c_{s}^{2}(1-i\tau_{eq}\omega)-2g_{1}(x)+8g_{2}(x)}+\frac{160(7-94g_{1}(x))}{8g_{1}(x)-33g_{2}(x)-2}.$ (72)
@@
$\omega_{1,2}=-\frac{i}{\tau_{eq}}\pm\frac{4k}{\sqrt{70}}\left(1-\frac{x^{2}}{24}+\cdots\right)+\mathcal{O}(k^{2}),$
$\omega_{3}=\frac{ix^{6}(\gamma_{E}-\ln{2})}{24\tau_{eq}}+ik^{2}\tau_{eq}\left(-\frac{1}{5}+\frac{x^{2}}{60}\right)+\mathcal{O}(k^{4}).$ (73)
The massless limit of the latter equation coincides with the analogous results in [27]. For large $x$, the corresponding retarded Green function reads as follows
$G^{0i,0i}(\omega,k)\bigg{|}_{x\gg 1}=\frac{(1-i\tau_{eq}\omega)\left(\varepsilon_{eq}+P_{eq}\right)\mathcal{N}^{(l)}_{shear}(x,\omega,k)}{\mathcal{D}^{(l)}_{shear}(x,\omega,k)},$ (74)
where
$\mathcal{N}^{(l)}_{shear}(x,\omega,k)=(1-i\tau_{eq}\omega)^{2}(h_{1}(x)-1)-\left(k^{2}\tau_{eq}^{2}+5(1-i\tau_{eq}\omega)^{2}\right)h_{2}(x),$
$\mathcal{D}^{(l)}_{shear}(x,\omega,k)=-\mathcal{N}^{(l)}_{shear}(x,\omega,k)-i\tau_{eq}\omega(1-i\tau_{eq}\omega)^{2}.$ (75)
The lowest lying modes are solutions of $\mathcal{D}^{(l)}_{shear}(x,\omega,k)=0$ which are
$\omega_{1,2}=-\frac{i}{\tau_{eq}}\pm\frac{ik}{x^{1/2}}+\frac{ik^{2}\tau_{eq}}{2x}+\mathcal{O}(k^{4}),$
$\omega_{3}=-\frac{ik^{2}\tau_{eq}}{x}+\mathcal{O}(k^{4}).$ (76)
All modes are stable when $x\to\infty$.
@@ V Analytical Behaviors
$\frac{1}{\omega}\mbox{Im}\,G_{00}^{(l)}(x;\omega,k=0)\bigg{|}_{\tau_{eq}T\to\infty}$ $=\pi\chi(x)\,\delta(\omega)\,\left(1+\frac{x^{2}K_{0}(x)+4xK_{1}(x)}{8K_{2}(x)}\right)$
$=\pi\chi_{x=0}\,\delta(\omega)\,\left(1-\frac{x^{2}}{64}\right),$ (78)
which is always positive and respects the positivity condition of the spectral density. We adopt $\chi_{x=0}=\chi(m=0)=\frac{T_{0}^{2}}{\pi^{2}}$. In the strong coupling regime, the expansion of the retarded two-point function matches Eq. (28), demonstrating that hydrodynamic poles emerge in this regime.
@@
This solution corresponds to the small-mass limit of $\omega{(h)}$ in Eq. (III.1) and coincides at $x=0$ with the results in [27]. In Fig. 4, we show the lowest-lying poles of the retarded $G_{00}$ correlator given in Eq. (28) for different $x$ values in the strong coupling regime at various $k\tau_{eq}$. Dashed lines indicate the positions of the branch cuts, and the hydrodynamic poles of the massless system are marked with crosses. Similar to Fig. 2, the mass tends to shift the poles toward the unstable region.
@@
$\frac{-1}{\omega}\mbox{Im}\,G^{00,00}(\omega,k=0)\bigg{|}_{\tau_{eq}T\to\infty}=\frac{\varepsilon_{eq}+P_{eq}}{c_{s}^{2}}\,\pi\,\delta(\omega)\,\left(1+g_{1}(x)-g_{2}(x)\right)=\frac{\varepsilon_{eq}+P_{eq}}{c_{s}^{2}}\,\pi\,\delta(\omega)\,\left(1+\frac{x^{2}}{24}\right),$ (81)
which satisfies the positive spectral density condition. In the strong coupling limit, the correlator expansion coincides with Eq. (65), and hydrodynamic poles naturally emerge. In Fig. 5, we depict the lowest-lying poles of $G^{00,00}$ for different $x$ values in the strong coupling regime at various $k\tau$. Unlike the massless case, all poles are located on the vertical line in the unstable region. These numerical solutions agree with those given in Eq. (IV.1). To analyze the results in the large $x$ limit, we present Figs. 6 and 7. In Fig. 6, we examine the pole structure of $G_{00}(x;\omega,k)$ for different values of $k\tau_{eq}$, each corresponding to poles at $x=(10,100,300)$. The insets precisely depict the pole locations for various $x$. As momentum increases, the poles shift toward positive imaginary values, indicating an instability at high momenta. Similarly, Fig. 7 illustrates the analytic structure of $G^{00,00}(\omega,k)$. Here, the poles appear in pairs with opposite real parts, characteristic of propagating modes.
Figure 4: Comparison between hydro poles for massless and massive cases in the charge correlation functions. In each figure, the dashed line indicates the location of the branch cut, and massless poles are shown with cross marks. Various ”$x$” ratios are displayed in different colors. Left (right) top plots correspond to $k\tau_{eq}=0.2(k\tau_{eq}=0.5)$, and left (right) bottom plots demonstrate $k\tau_{eq}=1(k\tau_{eq}=1.5)$.
@@
Figure 6: Sketch of poles for large massive particles in the charge correlation function. Massless poles are indicated with a cross sign. Inside plots show the precise location of poles for $x=(10,100,300)$.
Figure 7: Structure of poles for large massive particles $x=(10,100,300)$ in the sound channel. Insets realize the accurate positions of poles for different $x$.